### v3

# OpenTCM: A GraphRAG-Empowered LLM-based System for Traditional Chinese Medicine Knowledge Retrieval and Diagnosis

Abstract
Traditional Chinese Medicine (TCM) represents a rich repository of ancient medical knowledge that continues to play an important role in modern healthcare. Due to the complexity and breadth of the TCM literature, the integration of AI technologies is critical for its modernization and broader accessibility. However, this integration poses considerable challenges, including the interpretation of obscure classical Chinese texts and the modeling of intricate semantic relationships among TCM concepts. In this paper, we develop OpenTCM, an LLM-based system that combines a domain-specific TCM knowledge graph and Graph-based Retrieval-Augmented Generation (GraphRAG). First, we extract more than 3.73 million classical Chinese characters from 68 gynecological books in the Chinese Medical Classics Database, with the help of TCM and gynecology experts. Second, we construct a comprehensive multi-relational knowledge graph comprising more than 48,000 entities and 152,000 interrelationships, using customized prompts and Chinese-oriented LLMs such as DeepSeek and Kimi to ensure high-fidelity semantic understanding. Last, we integrate OpenTCM with this knowledge graph, enabling high-fidelity ingredient knowledge retrieval and diagnostic question-answering without model fine-tuning. Experimental evaluations demonstrate that OpenTCM achieves mean expert scores (MES) of 4.378 in ingredient information retrieval and 4.045 in diagnostic question-answering tasks, outperforming state-of-the-art solutions in real-world TCM use cases.
Index Terms:
Traditional Chinese Medicine, Large Language Model, Prompt, Knowledge Graph, GraphRAG
I Introduction
Traditional Chinese Medicine (TCM) represents a rich repository of ancient medical knowledge that continues to hold significant value in modern healthcare systems [1]. Due to the complexity of the literature and the practice of TCM, many researchers have tried to build systems to assist with consultation and diagnosis for TCM physicians. Current TCM knowledge retrieval and diagnosis systems predominantly often follow a paradigm of fine-tuning general-purpose Large Language Models (LLMs), such as ChatGPT [2] or Claude [3], on domain-specific datasets. However, they often face significant challenges such as the complexity of parsing classical Chinese texts, the intricate relationships among TCM concepts (such as ingredients, symptoms, and diagnostic principles), and the absence of structured knowledge such as knowledge graphs and retrieval-augmented generation (RAG) [4, 5]. In particular, ancient TCM literature is often written in archaic language, making them difficult for modern practitioners and computational models to interpret. Furthermore, existing approaches primarily focus on modern medical literature, overlooking the foundational corpus of classical TCM. Although recent advances in LLMs have proven to be effective in medical knowledge democratization, the direct application of general-purpose LLMs to TCM still faces significant limitations, such as hallucination in prescription generation and semantic understanding of classical Chinese.
To address these limitations, we propose OpenTCM, an LLM-based TCM knowledge retrieval and diagnosis system that integrates domain-specific knowledge with Retrieval-Augmented Generation with Graphs (GraphRAG) [6]. First, we extract over 3.73 million ancient Chinese characters from 68 gynecological books sourced from the Chinese Medical Classics Database [7], with verification from experienced TCM and gynecology experts. Next, we construct a comprehensive multi-relational knowledge graph containing over 48,000 entities and 152,000 interrelationships. This graph spans key TCM concepts, including 3,700+ ingredients, 14,000+ diseases, 17,000+ symptoms and treatments, and 65,000+ ingredient references. The construction process leverages customized prompts and domain-adapted LLMs, such as DeepSeek [8] and Kimi [9], to ensure accurate semantic understanding of classical Chinese medical literature. Compared to general-purpose prompts and LLMs, our approach achieves significant improvements in knowledge graph quality, with a precision of 98.55%, recall of 99.60%, F1-score of 99.55%, and accuracy of 98.17%. In addition, our experiments show that OpenTCM achieves a Mean Expert Score (MES), accuracy, and inter-rater agreement (IRA) of 4.378, 99.0%, and 0.057 for ingredient retrieval, and 4.045, 98.8%, -0.013 for diagnostic question-answering, respectively. By eliminating the need for model fine-tuning, OpenTCM leverages retrieval-augmented generation to provide accurate, context-aware responses with minimal computational overhead. Our structured knowledge graph also mitigates hallucination risks and enhances reasoning over complex TCM relationships, offering a scalable and efficient alternative to traditional methods.
Our main contributions are summarized as follows.
  • We develop OpenTCM, the first LLM-based TCM knowledge retrieval and diagnosis system that integrates a large corpus of structured TCM knowledge derived from ancient TCM literature, which consists of 68 gynecological books with over 3.73 million ancient Chinese characters from the Chinese Medical Classics Database.
  • We propose a TCM knowledge graph construction approach that combines customized prompts with domain-specific LLMs, achieving high-fidelity semantic extraction from classical texts.
  • We integrate the knowledge graph with GraphRAG to enhance reasoning capabilities over complex TCM interrelationships while maintaining computational efficiency.
  • Our experimental evaluation demonstrates that our system significantly outperforms existing solutions, achieving a precision of 98.55% in knowledge graph construction task, an MES of 4.378 and accuracy of 99.0% in ingredient retrieval, and an MES of 4.045 and accuracy of 98.8% in diagnostic question-answering.
The remainder of this paper is organized as follows: Section II reviews related work on TCM knowledge graphs and LLMs. Section III introduces our system architecture, including data collection, knowledge graph construction, and the GraphRAG framework. Section IV presents experimental evaluations. Finally, Section V outlines future work and Section VI concludes our work. Our project will be publicly available at https://github.com/XY1123-TCM/OpenTCM.
II Related Work
This section reviews the existing literature on TCM knowledge graphs, training-based LLMs for TCM, and the emerging fusion of knowledge graphs with LLMs via GraphRAG.
II-A TCM Knowledge Graphs
Knowledge graphs have emerged as a promising tool to structure TCM’s complex knowledge system. For example, Zhang et al. [10] utilize LLMs for TCM knowledge graph construction, employing named entity recognition and few-shot learning to reduce manual annotation efforts in applications such as education, diagnosis, and treatment. Similarly, Duan et al. [11] propose a hybrid method integrating LLMs with manual verification to build a TCM case knowledge graph, enabling a question-answering system. While these efforts demonstrate the potential of knowledge graphs in TCM, they do not incorporate the capacity to reason over relational complexities—a gap OpenTCM aims to address. In contrast, we incorporate GraphRAG in our OpenTCM system. enabling reasoning over relational complexities in TCM literature.
II-B LLMs for TCM
In recent years, driven by the success of LLMs in modern medical domains [12], the application of LLMs for TCM has attracted significant attention to enhance clinical reasoning and knowledge access. BianCang [13], built on the Qwen-2/2.5 architecture, undergoes continuous pre-training and fine-tuning to embed TCM expertise, improving its domain-specific performance. Lingdan [14], based on Baichuan2-13B-Base, employs a Chain-of-Thought approach, achieving an 18.39% improvement in Top20 F1-score for symptom analysis and prescription tasks compared to baseline solutions. Other models, such as Qibo [15] and Zhongjing [16], further showcase LLM’s potential in TCM consultations, with matching capabilities to ChatGPT with fewer parameters. TCMChat [17] uses pre-training and supervised fine-tuning on curated TCM text and Chinese question-answer datasets, focusing on tasks like entity extraction and ingredient recommendation. These studies, however, primarily rely on resource-intensive fine-tuning and lack structured knowledge integration, often resulting in computational inefficiency and hallucinations. In contrast, OpenTCM uses GraphRAG to realize high fidelity without computationally expensive pre-training or fine-tuning.
II-C Knowledge Graph-based Medical LLM
Recently, fusion of knowledge graphs with LLMs via GraphRAG has explored in medicine. The most related paper to our work is MedGraphRAG [18], which proposes a graph-based RAG framework for general medicine, outperforming state-of-the-art models across multiple benchmarks. However, its focus on modern medicine makes TCM-specific knowledge retrieval and diagnosis challenging. In this work, OpenTCM employs a training-free approach using GraphRAG, which not only reduces computational overhead but also integrates the rich relational data from the TCM literature to improve response accuracy and contextual relevance.
Figure 1: System architecture of OpenTCM, which consists of three main components: (1) data collection and pre-processing, (2) knowledge graph construction, and (3) graph-based retrieval-augmented generation.
III System Overview
In this section, we will describe the system overview of the OpenTCM. As shown in Figure 1, OpenTCM consists of three main components: (1) data collection and pre-processing, (2) knowledge graph construction, and (3) graph-based retrieval-augmented generation.
The process starts with collecting gynecological books from the Chinese Medical Classics database [7]. These medical literature, covering obstetrics, gynecology, and fertility, are carefully reviewed and corrected by TCM and gynecology medical experts to ensure data quality. Next, a multi-relational knowledge graph is constructed using well-customized prompts and advanced domain-adapted LLMs like Deepseek and Kimi. This knowledge graph captures key TCM elements such as ingredients, symptoms, and treatments, forming a structured representation of TCM. Last, the constructed knowledge graph serves as the backbone for the database, enabling the implementation of GraphRAG. This integration allows OpenTCM to offer powerful features such as TCM ingredient search and diagnostic question-answering without requiring pre-training or fine-tuning of the LLMs, significantly reducing computational overhead while maintaining high performance.
III-A Data Collection and Pre-processing
We extract 68 ancient classical books on traditional Chinese medicine in gynecology (20 Obstetrics, 43 Gynecology, and 5 Fertility) from the Digital Resource Library of Ancient Chinese Medicine Ancient Books of China, the most authoritative and systematic database of research on traditional Chinese medicine [7]. The books contain 6,787 chapters and 3,731,358 characters, covering diverse topics such as prescriptions, traditional Chinese patent medicines, simple preparations and medical case references (see Tables I and II). After collection, these books are then handed over to experienced TCM practitioners from Traditional Chinese Medicine and Gynecology for rigorous review, correction, and formatting to produce clean, usable dataset.
TABLE I: Number of books, chapters, and characters in different TCM specialties
Quantity Obstetrics Gynecology Fertility Total
Book 20 43 5 68
Chapter 1987 4496 304 6787
Character 734095 2813900 183363 3731358
TABLE II: Number of ingredients, diseases, symptoms, treatments and ingredient references in TCM Database
Ingredients Diseases Symptoms Treatments Ingredient-References
$37$ 14059 17031 17031 65847
III-B Knowledge Graph Construction
We use LLMs to extract structured knowledge graphs from the extensive corpus of ancient Chinese characters in TCM literature. However, there are two primary challenges: (1) accurately interpret the semantics of ancient Chinese texts, and then (2) elicit precise, structured responses—such as ingredients, diseases, symptoms, treatments and ingredient references. To overcome these issues, we leverage Chinese-oriented LLMs, which excel in capturing the linguistic and contextual nuances of Chinese literature, and employ carefully crafted prompts to steer the models toward generating detailed and accurate outputs. This methodology ensures the extraction of high-quality information, encompassing symptoms, prescriptions (e.g., ingredient names and dosages), and comprehensive treatment plans. Moreover, the construction of the knowledge graph relies on a systematic prompt-based extraction process tailored to TCM-specific data. Key components of our customized prompt design are as follows.
  • Role Definition: The prompt defines the model as a “TCM Data Processing Assistant,” emphasizing its proficiency in handling classical Chinese texts and its focus on extracting relevant information from TCM literature.
  • Task Description: The prompt instructs the model to extract structured information from the content of TCM chapters and output the results in JSON format. Specifically, the model is required to extract four types of information: (1) book chapter information, including chapter names, IDs, and content, which are stored in a “book” table; (2) treatment plan information, covering treatment names, indications, symptoms, and notes, stored in a “treatment” table; (3) ingredient information, including ingredient names and descriptions, stored in a “ingredient” table; and (4) ingredient-treatment relationship information, which involves ingredient names, dosages, and preparation methods, stored in a “ingredient_ref” table while ensuring the uniqueness of each combination of ingredient ID, treatment ID, and reference ID.
  • Structured Output: The prompt requires the model to output the extracted information in a pre-defined JSON format based on the specified table structures, facilitating subsequent knowledge graph construction and application.
  • Example Input and Output: The prompt includes a sample TCM text and the corresponding expected JSON output to illustrate the task requirements and guide the model to understand the desired extraction and formatting processes.
Through the integration of these models and customized prompts, we have successfully built a knowledge base comprising 3,737 ingredients, 17,031 treatments, and 65,847 ingredient references. The resulting TCM knowledge graph encapsulates 48,406 entities and 152,754 interrelationships, among which 10 types of triplets are formed, as depicted in Table III.
TABLE III: Constructed Interrelationships of TCM Knowledge Graph
Interrelationship Num Interrelationship Num
Belong to Category 48406 Include Section 294
Include Chapter 6786 Belong to book 6786
Treatment Plan 17001 Treat Disease 16133
Describe Disease 16104 Treatment Symptom 13605
Symptoms Present 13581 Ingredient Use 65846
The whole TCM Knowledge Graph and a partial graph are visualized in Figures 2 and 3. This structured knowledge graph not only preserves the richness of TCM data but also facilitates advanced analysis and application in modern research contexts.
Figure 2: Visualization of our TCM knowledge graph.
Figure 3: A partial visualization of our TCM knowledge graph.
TABLE IV: Performance of customized prompt verses general prompt, KIMI-based
General Customized
Precision 90.1% 98.55%
Recall 95.9% 99.60%
F1-score 92.3% 99.55%
Accuracy 86.8% 98.17%
III-C Graph-based Retrieval-Augmented Generation
We use GraphRAG [6, 19] to integrate graph-structured data into OpenTCM to enhance retrieval and generation processes. While conventional RAG approaches predominantly processes unstructured text or images, GraphRAG leverages the rich structural relationships embedded within a knowledge graph to improve both information access and reasoning capabilities. This is particularly valuable for TCM, which features complex inter-dependencies among ingredients, symptoms, treatments, and syndromes.
We adopt GraphRAG to operate effectively on the TCM knowledge graph by directly utilizing the multi-relational structure that is characteristic of traditional Chinese medicine theory. The knowledge graph includes over 48,000 entities and 152,000+ relationships such as treats, includes, associated with, and belongs to, which encode rich semantic information. By grounding retrieval in this structured graph rather than text, our system is able to more precisely trace clinically meaningful paths. For example, linking a symptom to a treatment through an associated ingredient or syndrome. This structure-aware approach allows the system to avoid irrelevant content and retrieve only contextually valid information, which is essential in domains where precise distinctions matter. Moreover, GraphRAG’s ability to traverse and aggregate information along multi-hop paths (e.g., symptom $\rightarrow$ syndrome $\rightarrow$ treatment $\rightarrow$ ingredient) enables OpenTCM to model the layered reasoning inherent in TCM practice. Instead of treating symptoms and ingredients as isolated keywords, GraphRAG interprets user queries within the relational graph, thereby capturing nuanced dependencies among concepts that are typically missed by flat retrieval systems.
These adaptations enhances our system’s performance by relying on a domain-specific graph structure. Therefore, OpenTCM mitigates hallucination and noise commonly observed in generic LLM outputs. This design supports high-precision tasks such as ingredient knowledge retrieval (accuracy: 99.0%) and diagnostic question-answering (accuracy: 98.8%), while reducing the need for model fine-tuning. The integration of GraphRAG with the TCM knowledge graph thus forms the core of OpenTCM’s ability to deliver accurate, interpretable, and context-aware responses in real-world medical scenarios.
Figure 4: An example case of OpenTCM diagnostic question-answering.
IV Experimental Evaluation
To evaluate OpenTCM, we conduct a comprehensive experimental evaluation focusing on three key capabilities: knowledge graph construction, TCM ingredient knowledge retrieval, and diagnostic question-answering.
IV-A Baseline Models
To ensure a thorough evaluation, we compare OpenTCM against several baseline models in both knowledge graph construction and real-world applications.
IV-A1 Knowledge Graph Construction Baselines
  • General-Purpose LLMs: We utilize ChatGPT-4 [2] and Claude-2 [20] as representatives of general-purpose large language models.
  • Domain-Specific LLMs: Deepseek [8] and Kimi [9] are employed as domain-adapted Chinese LLMs, fine-tuned on medical and classical Chinese corpora.
IV-A2 Real-World Application Baselines
  • General-Purpose LLMs: ChatGPT-4 [2] and Claude-2 [20] are used to assess the performance of general-purpose models in TCM applications.
  • TCM-Specific LLMs: BianCang [13] is a recent LLM specifically designed for TCM tasks, incorporating domain-specific knowledge and training data.
IV-B Knowledge Graph Construction
Evaluating the quality of the TCM knowledge graph is a cornerstone of this study, given its role as the foundation for OpenTCM’s downstream applications. In the absence of a pre-existing gold standard for TCM knowledge graph construction, we adopt a sampling inspection approach, enlisting 4 domain experts in TCM and Gynecology to assess the generated triples. Our evaluation examines two dimensions: (1) the effectiveness of customized versus general prompts, and (2) the performance of different backbone models, including Chinese-oriented LLMs (i.e., Deepseek [8], Kimi [9]) and more general LLMs (i.e., ChatGPT-4 [2], Claude [20]).
We task the LLMs with extracting 152,754 triples from a sampled subset of the corpus, comprising randomly selected chapters from the 68 gynecological books. In summary, the constructed TCM knowledge graph, containing over 48,000 entities and 152,000 relationships, is validated through expert review of randomly selected 600 chapters. To evaluate the knowledge graph of OpenTCM quantitatively, we randomly select 1795 triples with corresponding expert annotations.
IV-B1 Metric
This data enables a quantitative assessment using four standard metrics:
  • Precision: The proportion of generated triples that are correct, reflecting the accuracy of the extracted knowledge.
[…]
  • Recall: The proportion of actual correct triples successfully extracted by the LLM, indicating its coverage of the true knowledge.
[…]
  • F1-Score: The harmonic mean of Precision and Recall, providing a balanced measure of extraction quality.
    $PrecisionRecall\text{F1-Score}=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}F1-Score = 2$ ⋅ divide start_ARG Precision ⋅ Recall end_ARG start_ARG Precision + Recall end_ARG (3)
  • Accuracy: The proportion of correctly identified triples relative to the total ground truth, gauging overall correctness (though less commonly emphasized in generation tasks).
[…]
Here, TP (True Positives) denotes triples generated by the LLM and validated as correct by experts, FP (False Positives) represents incorrect triples generated by the LLM, and FN (False Negatives) indicates correct triples missed by the LLM.
IV-B2 Performance
We design specific prompts tailored to TCM data extraction, covering chapter information, treatment plans, ingredient details, and their relationships, which is more suitable to our task than the general-purpose prompt. The customized prompts demonstrate a clear advantage over general prompts (as shown in Table IV, KIMI-based), indicating their effectiveness in enhancing extraction accuracy and completeness. Besides, our tailored prompt is compatible with multiple LLM from different vendors, including Deepseek, Kimi, GPT4 and Claude2. We observe that DeepSeek and Kimi show superior performance as shown in Table V, we hypothesize that those models are more capable because they use more Chinese content as the training set. This significant difference highlights the importance of using models specifically adapted to the linguistic and contextual characteristics of TCM literature.
TABLE V: Performance Comparison among Different backbone LLMs in Knowledge Graph Construction
GPT4 Claude2 Deepseek KIMI
Precision 94.6% 94.26% 98.61% 98.55%
Recall 98% 97.41% 99.27% 99.60%
F1-score 95.6% 95.37% 98.49% 99.55%
Accuracy 92.8% 91.96% 97.9% 98.17%
An expert review of the full knowledge graph, encompassing over 48,000 entities and 152,000 relationships, confirms its fidelity in capturing TCM’s intricate knowledge structure. The combination of customized prompts and domain-adapted LLMs not only ensures high-quality triple extraction—as evidenced by the Precision, Recall, F1-Score, and Accuracy but also establishes a solid foundation for OpenTCM’s downstream applications like TCM ingredient search and diagnostic question-answering functionalities.
IV-C Real-world Applications
To assess OpenTCM’s effectiveness on real-world applications, we evaluate its performance on two critical downstream tasks: ingredient medicine information retrieval and diagnostic question answering. These tasks test the system’s ability to leverage the TCM knowledge graph and GraphRAG technology to provide accurate, contextually relevant, and user-friendly responses.
Due to the absence of a definitive ground truth, we collect and annotate a dataset consisting of 257 ingredient information search queries and 303 diagnostic consultation questions. Each query is processed by OpenTCM and several baseline systems, including general-purpose LLMs (KIMI) and recently proposed TCM-specific LLMs (BianCang). The system’s responses were evaluated by a panel of four medical experts, comprising two clinicians with more than 20 years of clinical experience and two medical Ph.D. students whose assessments were validated by their respective supervisors. A 5-point Likert scale, ranging from 1 (irrelevant) to 5 (highly relevant), was employed for scoring.
IV-C1 Metrics
We employ three key metrics for evaluating the performance of the system:
  • Mean Expert Score (MES): This metric represents the average score assigned by domain experts to assess the relevance and correctness of the retrieved or generated information. MES quantitatively evaluates the experts’ subjective judgment of the system’s response quality.
  • Accuracy: Accuracy is defined as the proportion of expert-labeled correct responses out of the total number of queries. For the purpose of this evaluation, a response is deemed correct if its expert score is equal to or greater than 3. This metric measures the overall capability of the model to produce acceptable responses.
  • Inter-Rater Agreement (IRA): IRA quantifies the consistency among multiple domain expert evaluators to ensure the reliability of the assessment results. For the calculation of IRA, expert ratings are binarized: scores greater than or equal to 3 are categorized as ”correct” (or ”agreement”), while scores less than 3 are categorized as ”incorrect” (or ”disagreement”). IRA reflects the objectivity of the evaluation by measuring the extent to which evaluators agree on these binary classifications.
IV-C2 Ingredient Information Retrieval
This task evaluates the model’s ability to return accurate and comprehensive information about ingredients such as “Danggui” and “Renshen”.
Experimental results are presented in Table VI. OpenTCM achieves the highest Mean Expert Score (4.378), surpassing both general-purpose models and TCM-specific baselines. This demonstrates the effectiveness of knowledge graph-enhanced reasoning in improving retrieval precision and contextual relevance.
Figure 5: Score distribution of different models for Ingredient Information Retrieval across four experts. OpenTCM shows significantly higher concentration at Score 4 and 5.
TABLE VI: Performance Comparison among OpenTCM and baseline solutions in Ingredient Knowledge Retrieval
Model MES Accuracy IRA
KIMI 2.691 53.9% -0.154
BianCang 2.616 47.1% -0.166
OpenTCM 4.378 99.0% 0.057
IV-C3 Diagnostic Question Answering
This task assesses the system’s diagnostic reasoning capabilities in clinical consultation scenarios (e.g., “How to treat menstrual disorders?” or “How to treat Qi deficiency with dampness?”). An example is illustrated in Figure 4.
Results in Table VII show that OpenTCM maintains strong performance across all metrics, particularly in interpretability and contextual coherence, outperforming general LLMs and remaining competitive with TCM-specific models.
TABLE VII: Performance Comparison among OpenTCM and baseline solutions in Diagnostic Question-Answering
Model MES ACC IRA
KIMI 3.043 50.2% -0.331
BianCang 2.455 44.7% -0.275
OpenTCM 4.045 98.8% -0.013
Figure 6: Score distribution of different models for Diagnostic Question Answering across four experts. OpenTCM responses receive notably higher expert ratings.
These results affirm OpenTCM’s practical value in real-world clinical and educational scenarios. Compared to both general-purpose and TCM-specific LLMs, OpenTCM exhibits superior performance in both factual accuracy and expert-perceived relevance, owing to its integration of structured domain knowledge through GraphRAG.
V Future Work
Future work will focus on expanding the TCM Knowledge Graph to include additional domains, such as rare diseases and historical case studies, to enhance OpenTCM’s applicability. We also aim to refine the GraphRAG framework to support more complex queries and integrate multimodal data, such as image recognition for raw and dried botanic and non-botanic substances. Additionally, developing tailored interfaces for practitioners, researchers, and patients will improve accessibility to TCM knowledge. Finally, we plan to collaborate with medical practitioners and institutions to integrate OpenTCM into healthcare systems such as in-home elderly care [21], clinical diagnostic [22], and exercise medicine [23, 24].
VI Conclusion
OpenTCM represents a significant advancement in applying LLMs to Traditional Chinese Medicine. By combining a meticulously constructed knowledge graph with GraphRAG, it achieves high accuracy and efficiency in knowledge retrieval and diagnostic tasks, paving the way for broader adoption of TCM practices.
Acknowledgment
This work is partly supported by Chinese Medicine Development Fund, Hong Kong SAR, China under 23B2/034A_R1.